Q3. (25 marks)Lower BoundExpected ValueUpper BoundSales quantity Sales price per unit Variable cost per unit Fixed cost95001000010500$9.75$10.00$10.25$4.80$5.20$5.60$15,000.00$18,000.00$21,000.00Initial requirement for equipment: S120,000 Depreciation: Straight-line to zero over the four-year life of the project with no salvage value. Required rate of return: 15% Marginal tax rate: 35%Calculate NPV according to best case scenario.
Answer #1

Answer:

Best case scenario will be when:

Sales quantity is at upper bound = 10,500 units

Sales price per unit is at upper bound = $10.25

Variable cost per unit is at lower bound = $4.80

Fixed cost is at lower bound = $15,000

Initial investment = $120,000

Annual depreciation = 120000 / 4 = $30,000

Annual Cash flows = ((Sales price - Variable cost per unit)* sales quantity - Fixed cost) * (1 - tax rate) + Depreciation * Tax rate

= ((10.25 - 4.80)* 10500 - 15000) *(1- 35%) + 30000 * 35%

= $37946.25

Best case NPV = Annual cash flow * PV factor for annuity - Initial Investment

= 37946.25 *((1 - 1 /(1+15%)^4)/15%) - 120000

= -11664.28

NPV according to best case scenario = -$11,664.28 or ($11,664.28)